Lauren plans to deposit $3000 into a bank account at the beginning of next month and $250/month into the same account at the end of that month and at the end of each subsequent month for the next 6 years. If her bank pays interest at a rate of 6%/year compounded monthly, how much will Lauren have in her account at the end of 6 years? (Assume she makes no withdrawals during the 6-year period. Round your answer to the nearest cent.)
$
Answer #1

Investment tenure(Nper)=6*12=72 months, Monthly rate=6%/12=0.5%, Initial Balance(PV)=3000, PMT=250, Future balance(FV)=?

ДА 1 PV 2 PMT 3 nper 4 rate 5FV -3000] -2500 72 0.50% =FV(B4,B3,B2,B1,0) FV(rate, nper, pmt, [pv], [type])

Α 1 PV 2 PMT 3 nper 4 rate -3000 -250 72 0.50% $25,898.35 5 FV 6

Hence, she will have $25898.35 at the end of 6 year period.